OpenStudy (anonymous):

Please check, I got 61. @jdoe0001

OpenStudy (anonymous):

I did too

OpenStudy (jdoe0001):

wonder what the common ratio is =)

OpenStudy (anonymous):

You got 61? :)

OpenStudy (jdoe0001):

hmmm nope... not yet.... did you find a common ratio?

OpenStudy (anonymous):

Yeah. at 1st=1 at 2nd=1+5=6 at 3rd=6+7=13 at 4th=13+9=22 at 5th=22+11=33 at 6th=33+13=46 at 7th=46+15=61

OpenStudy (jdoe0001):

hmmmm so.. what's the pattern then?

OpenStudy (anonymous):

Idk. 2?

OpenStudy (anonymous):

Is 61 correct as my final answer?

OpenStudy (jdoe0001):

..... anyhow... I guess you can use 61 based on that so what you got is just \(\bf \large n^2+2(n-1)\)

OpenStudy (jdoe0001):

so, try 7 there

OpenStudy (anonymous):

61, *ding, ding, ding!*

OpenStudy (jdoe0001):

heh yes